# Online Budget-Feasible Mechanism Design with Predictions

### v1

@@ Abstract
Augmenting the input of algorithms with predictions is an algorithm design paradigm that suggests leveraging a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is perfect (consistency), while also providing a performance guarantee when the prediction fails (robustness). Recently, xu2022mechanism and agrawal2022learning proposed to consider settings with strategic agents under this framework. In this paper, we initiate the study of budget-feasible mechanism design with predictions. These mechanisms model a procurement auction scenario in which an auctioneer (buyer) with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, so as to maximize her own valuation function. We focus on the online version of the problem where the arrival order of agents is random. We design mechanisms that are truthful, budget-feasible, and achieve a significantly improved competitive ratio for both monotone and non-monotone submodular valuation functions compared to their state-of-the-art counterparts without predictions. Our results assume access to a prediction for the value of the optimal solution to the offline problem. We complement our positive results by showing that for the offline version of the problem, access to predictions is mostly ineffective in improving approximation guarantees.
@@ 1 Introduction
Our work evolves around the design of auction mechanisms for procurement auctions. Procurement auctions, also known as reverse auctions, are mechanisms for assigning a set of projects or tasks to a set of candidate providers. In the more traditional applications of reverse auctions, the auctioneer may be a governmental organization (e.g., for assigning a public project to a construction firm), or a company interested in outsourcing tasks, such as the supply of logistics or retail services; see CSS06. In recent years, the explosion and monetization of the Internet, have created even further applications and platforms where even a single individual can conduct online assignments/auctions. Two such prominent families of examples include platforms for online labour markets, such as Upwork or Freelancer, and platforms for crowdsourcing systems, where the posted tasks can range from very elementary assignments as image labeling, all the way to more specialized services like translating documents. Participatory crowdsensing also falls under this framework, where some entities are interested in collecting sensing information from smartphones or other devices.
@@
The emergence of such online markets, has caught the increasing attention of the research community, which has focused on properly formulating and addressing relevant research questions. One of the popular models in the literature for capturing procurement auctions is that of budget-feasible mechanism design, originally proposed in the seminal paper of singer2010budget. This model concerns a scenario where an auctioneer with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, who may misreport their cost to their advantage for obtaining higher payments. Under these considerations, a natural goal for the auctioneer is to come up with a truthful mechanism for hiring a subset of the agents that maximizes the procured value (measured by the valuation function of the auctioneer), and such that the total payments to the agents respect the budget limitation. Given that even the non-strategic versions of such budget-constrained problems tend to be NP-hard, the main focus is on providing budget-feasible mechanisms that achieve approximation guarantees on the auctioneer’s optimal potential value. This has led to a steady stream of works which we outline in our related work section.
@@
The focus of our work is to study budget-feasible mechanisms under the rather recently introduced framework of learning-augmented mechanism design. This forms a new paradigm in designing algorithms, where the input is augmented by a prediction on some relevant parameter of the problem, possibly obtained by past data. The main goal is to leverage a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is accurate (referred to as consistency), while also providing a performance guarantee when it is not (referred to as robustness), without knowing the prediction accuracy in advance. Recently, xu2022mechanism and agrawal2022learning proposed to consider settings with strategic agents through this lens. Motivated by these two initial attempts, there has been a fast growing interest and a series of other works, that studied the design of truthful mechanisms with predictions in a variety of settings, leading mostly to positive results (i.e., improved approximation guarantees under correct predictions while not far off from the best known guarantee otherwise).
Our Contribution.
@@
  • In Section 3 we propose a family of universally truthful, budget-feasible mechanisms (LABEL:mechanism:mono-sm-conv; parameterized by an error tolerance parameter $\tau$) for monotone submodular objectives which attain $O(1)$ consistency and robustness. For $\tau$ close to $1$, the consistency, i.e., the approximation guarantee when the prediction is perfect, can be as low as $6$, which is a significant improvement even versus the best known approximation guarantee of $54.4$ for Submodular Knapsack Secretary, i.e., the non-strategic version of the problem, without predictions (due to feldman11).
  • For smaller values of $\tau$, the robustness of LABEL:mechanism:mono-sm-conv in Section 3, i.e., the approximation guarantee when the prediction is arbitrarily bad, can be as low as $146$. This also implies a $146$-approximation mechanism without predictions, improving the previously known approximation of $1710$ by amanatidis2022budget by a factor greater than $11$!
  • In LABEL:sec:mono-sm-sampling, we present a different mechanism (LABEL:mechanism:mono-sm2) where the prediction implicitly determins the length of the initial agent-sampling window (although here we have fixed the latter for the sake of presentation). Compared to LABEL:mechanism:mono-sm-conv, this approach is more conservative, as LABEL:mechanism:mono-sm2 is, in a sense, less sensitive to (very) erroneous predictions. In particular, we show that our mechanism achieves a consistency-robustness tradeoff of $95$ and $280$.
  • We obtain analogous results for non-monotone submodular objectives, albeit with some loss on approximation, via LABEL:mechanism:nonmono-sm-conv and LABEL:mechanism:non-mono-sm2. Specifically, for $\tau$ close to $1$, the consistency of LABEL:mechanism:nonmono-sm-conv can be as low as $19$, whereas, for smaller values of $\tau$, its robustness can be as low as $445$. We emphasize once more that, even when one completely ignores the prediction, our analysis leads to a large improvement on the current state-of-the-art approximation due to amanatidis2022budget, which is $1710$ even for Submodular Knapsack Secretary (the non-strategic version of the problem). Finally, we show that the result of the analogue of LABEL:mechanism:mono-sm2 for non-monotone submodular valuations (LABEL:mechanism:non-mono-sm2) achieves a consistency-robustness tradeoff of $228$ and $818$.
  • We complement our positive results by showing in LABEL:sec:lower-bound that for the offline version, access to predictions is mostly ineffective in improving approximation guarantees. In particular, we show that no randomized, universally truthful, budget-feasible mechanism with bounded robustness can achieve a consistency guarantee better than the lower bound of $2$ that is known to hold for additive valuations and randomized mechanisms (chen2011approximability).
@@
The design of truthful budget-feasible mechanisms was initiated by singer2010budget for additive and monotone submodular valuation functions, and has sparked a rich line of work. For additive valuation functions a best-possible $2$-approximation randomized mechanism was given by gravin2020optimal, whereas it is also known that the best possible factor of any deterministic mechanism is between $1+\sqrt{2}$ and $3$ (due to chen2011approximability and gravin2020optimal). For monotone submodular valuation functions, through a series of improvements (see e.g., chen2011approximability; jalaly2021simple; balkanski2022deterministic), the current state of the art is $4.45$ for deterministic and $4.08$ for randomized mechanisms due to han2024triple who follow the clock auction paradigm. Without the monotonicity assumption, obtaining any $O(1)$ approximation seemed to be more challenging (see e.g., amanatidis2022budget), but here as well the recent improvements of balkanski2022deterministic; huang2023randomized have led to a relatively small approximation ratio, namely $11.67$, again by han2024triple.
@@
There have also been works focusing on richer valuation classes like XOS and subadditive functions (see e.g. bei2017worst; amanatidis2017budget; neogi_et_al:LIPIcs.ITCS.2024.84), exploring beyond-the-worst-case scenarios (rubinstein23), or using large market assumptions, see AnariGN18; jalaly2021simple. We refer the reader to the recent survey of liu24 for an overview of various settings (e.g., environments with combinatorial constraints and more general single-parameter domains).
@@
A recent direction in theoretical computer science is the area of “beyond the worst-case analysis” (see, e.g., roughgarden2021beyond). One prominent approach within this framework is algorithms with predictions, which originated in the field of online algorithms (see, e.g., lykouris2021competitive) and utilizes predictions, e.g., provided by an ML algorithm trained on historical data, to overcome worst case lower bounds. The idea has been also applied in strategic settings, where predictions are used to simulate (an aggregation of) the agents’ private information, starting with the work of agrawal2022learning and xu2022mechanism. Areas of application in mechanism design include auction-related environments (see e.g., prasad23; balkanski23; christodoulou24; caragiannis24; balkanski23scheduling; gkatzelis25), online mechanisms (see e.g., balkanski23; lu24), mechanisms without money (see e.g. balkanski24randomized; colini24; barak24; cohen24) and computational social choice (see e.g. filos2025). For an almost exhaustive list of papers in the area, see the online repository of algorithms_with_predictions.
@@
The online budget-feasible mechanism design problem we study was introduced by badanidiyuru2012learning and is a generalization to the strategic setting of the Submodular Knapsack Secretary problem (see, e.g., babaioff17). The latter has also been considered by bateni13 and subsequently by feldman11 and KesselheimT17. Very recently, charalampopoulos25 devised a truthful mechanism which achieves a constant approximation (for a very large constant) for posted pricing and monotone submodular valuations. The state-of-the-art approximation for monotone or non-monotone submodular valuation functions for the strategic version we study here is due to amanatidis2022budget, who achieve a competitive ratio of $1710$. Note that the Secretary problem (additive objective and cardinality constraint) has been studied in environments with predictions (see, e.g., antoniadis23; perchet23; fujii24; balkanski24fair), yet our work is the first to study the online budget-feasible mechanism design problem—and its algorithmic counterpart—in such augmented environments.
@@ 2 Preliminaries
  • individually rational; no agent can be hired for less than they asked for, i.e., $p_{i}(\mathbf{b})\geq b_{i}$, if $i\in A(\mathbf{b})$
@@
$\max\ v(S)\quad\text{s.t.}\quad\textstyle\sum_{i\in S}c_{i}\leq B,\quad S\subseteq N.\vspace{-3pt}$ (1)
@@ 3 A Natural Prediction-Augmented Mechanism
In this section, we assume that the valuation function $v$ is monotone submodular. We address the more general, non-monotone case in LABEL:sec:nonmono-submodular.

### v3

@@ Abstract
Augmenting the input of algorithms with predictions is an algorithm design paradigm that suggests leveraging a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is perfect (consistency), while also providing a performance guarantee when the prediction fails (robustness). Recently, Xu and Lu (2022) and Agrawal et al. (2024) proposed to consider settings with strategic agents under this framework. In this paper, we initiate the study of budget-feasible mechanism design with predictions. These mechanisms model a procurement auction scenario in which an auctioneer (buyer) with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, so as to maximize her own valuation function. We focus on the online version of the problem where the arrival order of agents is random. We design mechanisms that are truthful, budget-feasible, and achieve a significantly improved competitive ratio for both monotone and non-monotone submodular valuation functions compared to their state-of-the-art counterparts without predictions. Our results assume access to a prediction for the value of the optimal solution to the offline problem. We complement our positive results by showing that for the offline version of the problem, access to predictions is mostly ineffective in improving approximation guarantees.
@@ 1 Introduction
Our work evolves around the design of auction mechanisms for procurement auctions. Procurement auctions, also known as reverse auctions, are mechanisms for assigning a set of projects or tasks to a set of candidate providers. In the more traditional applications of reverse auctions, the auctioneer may be a governmental organization (e.g., for assigning a public project to a construction firm), or a company interested in outsourcing tasks, such as the supply of logistics or retail services; see Cramton et al. (2006). In recent years, the explosion and monetization of the Internet, have created even further applications and platforms where even a single individual can conduct online assignments/auctions. Two such prominent families of examples include platforms for online labour markets, such as Upwork or Freelancer, and platforms for crowdsourcing systems, where the posted tasks can range from very elementary assignments as image labeling, all the way to more specialized services like translating documents. Participatory crowdsensing also falls under this framework, where some entities are interested in collecting sensing information from smartphones or other devices.
@@
The emergence of such online markets, has caught the increasing attention of the research community, which has focused on properly formulating and addressing relevant research questions. One of the popular models in the literature for capturing procurement auctions is that of budget-feasible mechanism design, originally proposed in the seminal paper of Singer (2010). This model concerns a scenario where an auctioneer with a strict budget constraint seeks to purchase goods or services from a set of strategic agents, who may misreport their cost to their advantage for obtaining higher payments. Under these considerations, a natural goal for the auctioneer is to come up with a truthful mechanism for hiring a subset of the agents that maximizes the procured value (measured by the valuation function of the auctioneer), and such that the total payments to the agents respect the budget limitation. Given that even the non-strategic versions of such budget-constrained problems tend to be NP-hard, the main focus is on providing budget-feasible mechanisms that achieve approximation guarantees on the auctioneer’s optimal potential value. This has led to a steady stream of works which we outline in our related work section.
@@
The focus of our work is to study budget-feasible mechanisms under the rather recently introduced framework of learning-augmented mechanism design. This forms a new paradigm in designing algorithms, where the input is augmented by a prediction on some relevant parameter of the problem, possibly obtained by past data. The main goal is to leverage a (possibly erroneous) prediction to improve worst-case performance guarantees when the prediction is accurate (referred to as consistency), while also providing a performance guarantee when it is not (referred to as robustness), without knowing the prediction accuracy in advance. Recently, Xu and Lu (2022) and Agrawal et al. (2024) proposed to consider settings with strategic agents through this lens. Motivated by these two initial attempts, there has been a fast growing interest and a series of other works, that studied the design of truthful mechanisms with predictions in a variety of settings, leading mostly to positive results (i.e., improved approximation guarantees under correct predictions while not far off from the best known guarantee otherwise).
Our Contribution
@@
  • In Section 3 we propose a family of universally truthful, budget-feasible mechanisms (Mechanism 1; parameterized by an error tolerance parameter $\tau$) for monotone submodular objectives which attain $O(1)$ consistency and robustness. For $\tau$ close to $1$, the consistency, i.e., the approximation guarantee when the prediction is perfect, can be as low as $6$, which is a significant improvement even versus the best known approximation guarantee of $54.4$ for Submodular Knapsack Secretary, i.e., the non-strategic version of the problem, without predictions (due to Feldman et al. (2011)).
  • For smaller values of $\tau$, the robustness of Mechanism 1 in Section 3, i.e., the approximation guarantee when the prediction is arbitrarily bad, can be as low as $146$. This also implies a $146$-approximation mechanism without predictions, improving the previously known approximation of $1710$ by Amanatidis et al. (2022) by a factor greater than 1111!
  • In Section 4, we present a different mechanism (Mechanism 4) where the prediction implicitly determines the length of the initial agent-sampling window (although here we have fixed the latter for the sake of presentation). Compared to Mechanism 1, this approach is more conservative, as Mechanism 4 is, in a sense, less sensitive to (very) erroneous predictions. In particular, we show that our mechanism achieves a consistency-robustness tradeoff of $95$ and $280$.
  • We obtain analogous results for non-monotone submodular objectives, albeit with some loss on approximation, via Mechanism 5 and Mechanism 8. Specifically, for $\tau$ close to $1$, the consistency of Mechanism 5 can be as low as $19$, whereas, for smaller values of $\tau$, its robustness can be as low as $445$. We emphasize once more that, even when one completely ignores the prediction, our analysis leads to a large improvement on the current state-of-the-art approximation due to Amanatidis et al. (2022), which is $1710$ even for Submodular Knapsack Secretary (the non-strategic version of the problem). Finally, we show that the result of the analogue of Mechanism 4 for non-monotone submodular valuations (Mechanism 8) achieves a consistency-robustness tradeoff of $228$ and $818$.
  • We complement our positive results by showing in Section 6 that for the offline version, access to predictions is mostly ineffective in improving approximation guarantees. In particular, we show that no randomized, universally truthful, budget-feasible mechanism with bounded robustness can achieve a consistency guarantee better than the lower bound of $2$ that is known to hold for additive valuations and randomized mechanisms (Chen et al., 2011).
@@
The design of truthful budget-feasible mechanisms was initiated by Singer (2010) for additive and monotone submodular valuation functions, and has sparked a rich line of work. For additive valuation functions a best-possible $2$-approximation randomized mechanism was given by Gravin et al. (2020), whereas it is also known that the best possible factor of any deterministic mechanism is between $1+\sqrt{2}$ and $3$ (due to Chen et al. (2011) and Gravin et al. (2020)). For monotone submodular valuation functions, through a series of improvements (see e.g., Chen et al. (2011); Jalaly and Tardos (2021); Balkanski et al. (2025)), the current state of the art is $4.45$ for deterministic and $4.08$ for randomized mechanisms due to Han et al. (2025) who follow the clock auction paradigm. Without the monotonicity assumption, obtaining any $O(1)$ approximation seemed to be more challenging (see e.g., Amanatidis et al. (2022)), but here as well the recent improvements of Balkanski et al. (2025); Huang et al. (2023) have led to a relatively small approximation ratio, namely $11.67$, again by Han et al. (2025).
@@
There have also been works focusing on richer valuation classes like XOS and subadditive functions (see e.g. Bei et al. (2017); Amanatidis et al. (2017); Neogi et al. (2024, 2025a)), exploring beyond-the-worst-case scenarios (Rubinstein and Zhao (2023)), or using large market assumptions, see Anari et al. (2018); Jalaly and Tardos (2021). Very recently the problem has been considered in a multi-parameter domain by Neogi et al. (2025b). We refer the reader to the recent survey of Liu et al. (2024) for an overview of various settings (e.g., environments with combinatorial constraints and more general single-parameter domains).
@@
A recent direction in theoretical computer science is the area of “beyond the worst-case analysis” (see, e.g., Roughgarden (2021)). One prominent approach within this framework is algorithms with predictions, which originated in the field of online algorithms (see, e.g., Lykouris and Vassilvitskii (2021)) and utilizes predictions, e.g., provided by an ML algorithm trained on historical data, to overcome worst case lower bounds. The idea has been also applied in strategic settings, where predictions are used to simulate (an aggregation of) the agents’ private information, starting with the work of Agrawal et al. (2024) and Xu and Lu (2022). Areas of application in mechanism design include auction-related environments (see e.g., Prasad et al. (2023); Balkanski et al. (2024b); Christodoulou et al. (2024); Caragiannis and Kalantzis (2024); Balkanski et al. (2023); Gkatzelis et al. (2025)), online mechanisms (see e.g., Balkanski et al. (2024b); Lu et al. (2024)), mechanisms without money (see e.g. Balkanski et al. (2024a); Colini-Baldeschi et al. (2024); Barak et al. (2024); Cohen et al. (2024)) and computational social choice (see e.g. Filos-Ratsikas et al. (2025)). For an almost exhaustive list of papers in the area, see the online repository of Lindermayr and Megow (2025).
@@
The online budget-feasible mechanism design problem we study was introduced by Badanidiyuru et al. (2012) and is a generalization to the strategic setting of the Submodular Knapsack Secretary problem (see, e.g., Babaioff et al. (2007)). The latter has also been considered by Bateni et al. (2013) and subsequently by Feldman et al. (2011) and Kesselheim and Tönnis (2017). Very recently, Charalampopoulos et al. (2025) devised a truthful mechanism which achieves a constant approximation (for a very large constant) for posted pricing and monotone submodular valuations. The state-of-the-art approximation for monotone or non-monotone submodular valuation functions for the strategic version we study here is due to Amanatidis et al. (2022), who achieve a competitive ratio of $1710$. Note that the Secretary problem (additive objective and cardinality constraint) has been studied in environments with predictions (see, e.g., Antoniadis et al. (2023); Benomar and Perchet (2023); Fujii and Yoshida (2024); Balkanski et al. (2024c)), yet our work is the first to study the online budget-feasible mechanism design problem—and its algorithmic counterpart—in such augmented environments.
@@ 2 Preliminaries
  • individually rational; no agent can be hired for less than they asked for, i.e., if $i\in A(\mathbf{b})$, then $p_{i}(\mathbf{b})\geq b_{i}$.
@@
$\max\ v(S)\quad\text{s.t.}\quad\textstyle\sum_{i\in S}c_{i}\leq B,\quad S\subseteq N.\vskip-3.0pt$ (1)
@@ 3 A Natural Prediction-Augmented Mechanism
In this section, we assume that the valuation function $v$ is monotone submodular. We address the more general, non-monotone case in Section 5.